Date:
Friday, October 23, 2020
Telescope:
CTIO 1.5-m
Submitter:
Roberto Aviles
Submitter Email:
Observer Support:
None
Night Assistant:
None
Problems:
None.
Program:
Proposal:
SMARTS K Kids
Target of Oportunity:
No
Time Observed:
9.75
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
9.75
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2020-10-23 | 9.75 | 0.0 | 0.0 | 0.0 | 0.0 | 9.75 |
For the record:
In the Afternoon: The fiber was aligned in the coude room end as well as at the Front End side (in the GAM) .
The pick-off mirror was cleaned using CO2.
We reported to Marco B. the Pfeiffer vacuum meter is not automatically turning OFF, situation that will be corrected tomorrow, By now it was manually turned OFF.
In the Night: The pointing, focus and tracking has been ok all night long.
The guider box was verified and found current position is (291,249.)
Two RV standards were repeatedly observed in slicer mode.
One standard for flux calibration was observed.
Plus, a Telluric standard.
All the possible decker modes were used and also, one standard was observed in different exposure times (HIP008102 with exposure times of 10, 60, 60 + ThAr Lamp + 180, 240, 360 + ThAr lamp + 450, 600 seconds + ThAr lamp.
Underline the original program was modified on demand and at the end we also added HIP029611 and HIP034222, 900 seconds each in slicer mode plus ThAr lamps.
Note: HIP034222 looks dim, we used a 100ms time for guiding.
Note: Leonardo Paredes supported observations, all night long!